OpenStudy (anonymous):

Find the standard form of the equation of the parabola with the given characteristics. Vertex (3,-9) ; Focus (3,-7)

OpenStudy (anonymous):

this may help... http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php

OpenStudy (anonymous):

Thanks! The equation for the standard form of a parabola would be y=ax^2+bx+c, right? How do I incorporate the vertex and focus into that formula?

OpenStudy (anonymous):

you have the vertex V(h,k) right?

OpenStudy (anonymous):

Yes, h=3, k=-9

OpenStudy (anonymous):

y-k = a(x-h)^2 for the vertex form at standard form y = ax^2 + bx + c x-h = a(y-k)^2 for sideways wherein x=ay^2 + by + c

OpenStudy (anonymous):

from the given points, the focus is 2 units above the vertex and both 3 units from x=0 line or simply the y-axis... thru these observations, we have a parabola which vertex and focus points located in the 4th quadrant and opens upward... :)

OpenStudy (anonymous):

so we can use the vertex form y-k = a (x-h)^2

OpenStudy (anonymous):

a = 1/(4p) where p is the distance between the vertex and focus points

OpenStudy (anonymous):

for a parabola that opens upward, a > 0...

OpenStudy (anonymous):

substitute all known values to vertex form then work on the equation until you got the standard form of Quadratic Equation....

OpenStudy (anonymous):

Oh okay! Thanks so much. And to be sure, p would equal 2 in this case because that's the distance right?

OpenStudy (anonymous):

yup... you're right... :)

OpenStudy (anonymous):

I'm at this point so far: y+9 = 1/8(x-3)^2. Am I on the right track?

OpenStudy (anonymous):

yup you can continue...

OpenStudy (anonymous):

Ty :)

OpenStudy (anonymous):

I'm not sure if this is right but I got: 1/8x^2 - 3/4x+81/8

OpenStudy (anonymous):

hmmm... there is something wrong with your simplification....

OpenStudy (anonymous):

you have to work on both side of the equation \[y+9=\frac{ 1 }{ 8 }(x-3)^2\]

OpenStudy (anonymous):

Oh okay. So would we expand (x-3)^2 first?

OpenStudy (anonymous):

we can start on that...

OpenStudy (anonymous):

(x-3)(x-3) would be x^2-6x+9 right?

OpenStudy (anonymous):

correct...

OpenStudy (anonymous):

And then you multiply that by 1/8 and get: 1/8x^2 -3/4+9/8

OpenStudy (anonymous):

you can have that, but much better we avoid fraction distribution, so to avoid doing how will we removed 1/8 on the right-hand side?

OpenStudy (anonymous):

multiplying everything by 8? i think.

OpenStudy (anonymous):

ok do it...

OpenStudy (anonymous):

I have 8y=x^2-6x-63 so far.

OpenStudy (anonymous):

now to follow the standard form of y=ax^2+bx+c...what must you do?

OpenStudy (anonymous):

Divide both sides by 8 right? to isolate the y.

OpenStudy (anonymous):

yup... then simplify further :)

OpenStudy (anonymous):

\[y=x^2/8 - 6x/8 - 63/8\]

OpenStudy (anonymous):

Hmmm, I'm not sure how to fix this haha

OpenStudy (anonymous):

well it's all that we need... merely the simplified one... \[y=\frac{ 1 }{ 8 }x^2-\frac{ 3 }{ 4 }x-\frac{ 63 }{ 8 }\]

OpenStudy (anonymous):

Oh so thats the simplified answer?

OpenStudy (anonymous):

congratulations you made it all by yourself.... :)

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

no problem... anytime... :)